# USPilot: An Embodied Robotic Assistant Ultrasound System with Large Language Model Enhanced Graph Planner

### v1

@@ III-D Ultrasound knowledge embedded and user intent understanding
Ours LLaMA3 8b 97.29 88.87 78.64 89.37 74.35 54.6 91.09 74.94 59.6
IV Experiment
In this section, we present several experiments designed to assess the performance of the USPilot’s components. The LLMEG module was evaluated on a publicly available dataset compared to other methods. The semantic router and ultrasound knowledge adapter were also validated on a pre-collected dataset. Furthermore, the USPilot system was integrated with our automated robotic ultrasound scanning system, demonstrating its functionality in a physical environment.
IV-A Comparison Experiment for LLMEG
We evaluated the proposed LLMEG with three public datasets from TaskBench[28] for instructed API selection or task planning. Multimedia dataset: This dataset includes user-centric tasks such as file downloading and video editing. It is represented as a graph with 40 vertexes and 449 edges. HuggingFace dataset: This dataset contains various AI models covering tasks across language, vision, audio, and video domains. Each task is treated as a tool for addressing specific problems, forming a graph with 23 vertexes and 225 edges. Dailylife APIs: This dataset comprises tools for daily services such as web search and shopping. It is represented as a graph with 40 vertexes and 1,560 edges.
We employed 3 metrics to evaluate the methods’ performance. Accuracy is the successful rate of completing the given tasks. The vertex F1 score evaluates the labels assigned to individual vertexes, which can be indicated as
, where the $Precision_{vertex}$ is the number of correctly identified vertexes among all predicted vertexes. $Recall_{vertex}$ is the number of correctly identified vertexes out of the total actual vertexes, representing correct API calls. The edge F1 measures the model’s ability to predict valid or expected API call sequences, which can be assigned as
, where $Precision_{edge}$ is the number of correctly identified edges among all predicted edges, and $Recall_{edge}$ is the number of correctly identified edges out of the total actual edges in API call sequences.
@@ IV-A Comparison Experiment for LLMEG
In the compression experiment, we employed two alternative methods: LLM+LLM: The LLM+LLM method involves collaboration between two LLMs. Initially, the text description of all the tools and the user request is provided to the first LLM, which generates the necessary substeps to complete the task. Subsequently, the generated substeps are combined with the tool descriptions to form another prompt for the second LLM, which selects the tools to be utilized. The selected tools are considered as vertexes, while two adjacent vertexes can be regarded as edges. GNN4Task: The GNN4Task generates subtasks using LLMs and employs a small language model (e5-355M) to embed each subtask. A GNN processes each embedded subtask to identify vertexes, with edges defined as connections between adjacent vertexes.
@@
We employed the 8-billion-parameter LLaMA3 model in our proposed method LLMEG and compared it with GNN4Task across the Dailylife, Multimedia, and HuggingFace datasets. Notably, GNN4task is equipped with one of the powerful LLMs, GPT4-tubo, a mixture of eight 222 billion parameters experts. As shown in Table I, the proposed method LLMEG achieves superior performance across all three datasets compared with the GNN4Task method. In the Dailylife dataset, LLMEG achieves vertex and edge F1 scores of 97.29% and 88.87%, respectively, outperforming GNN4Task (97.10% and 88.52%). Both methods perform well on this dataset due to the inherent ability of LLMs to understand daily service descriptions. For the Multimedia dataset, our method improves vertex and edge F1 scores by 0.81% and 4.75%, respectively, over GNN4Task. On the more professional HuggingFace dataset, our proposed method significantly enhances vertex selection and edge prediction by 14.8% and 24.45%, respectively. We also compared our method with a variant that replaces the LLM-generated subgraph with depth-first search (DFS) for planning. Using subgraph planning, the edge F1 score increases by 62%, 23.4%, and 22.4%, respectively, while the total accuracy increases by 39.52%, 26.6%, and 25.8% in the datasets. These results highlight the importance of semantic understanding in generating correct API execution sequences, even when API selection is accurate. The improvement in edge prediction varies between datasets and is particularly pronounced in the Dailylife dataset, which contains an average of 39 edges per vertex. As graph complexity increases, the significance of semantic information becomes more evident.
@@
After employing the LLMEG, we conducted experiments using different LLMs with varying parameter sizes to assess planning capabilities. As shown in Fig. 5(a), our proposed method outperforms the original LLM-based method and GNN4Task in both vertex and edge F1 scores by over 40% across all datasets when using LLaMA3-8b and its embeddings. For other LLMs, LLMEG also abstracts word embedding with 137M parameters small language model (nomic-embed-textv1.5). Our proposed method achieves better performance across all metrics for the similar size LLM Qwen2.5-7b. However, when integrated with smaller LLMs shown in Fig. 5(c) and Fig. 5(d), our LLMEG can result a high value only in vertex F1 score but struggles with edge F1 scores and accuracy. This is due to smaller LLMs’ limited semantic understanding capabilities, which are insufficient for effective subgraph replanning.
IV-B User intent understanding and ultrasound knowledge embedded
@@
The user intent prediction MLP was trained with 1.2K instructions on robot tasks, general question-answering (QA), and ultrasound-related questions with an equal ratio of 1:1:1. We used sigmoid as the activation function to classify the user inputs as QA inquiries or ultrasound robot commands. As shown in Table II, the model was evaluated on a 14K dataset, comprising 9K QA instances and 5K robotic commands. The results demonstrate the MLP classification layer can reach an accuracy of 100% for QA intent and 97% for robotic command intent.
In addition, the ultrasound robotic adapter was fine-tuned using 2K instructions on seven body parts: carotid vessel, thyroid, liver, kidney, spleen, femoral artery, and gallbladder. The dataset comprises instructions and expected subtasks, in which the instructions have the patient’s direct command and symptom description, such as “Scan the patient’s liver” and “The patient has jaundice; check him with ultrasound.” An evaluation set was prepared for six body parts and functional commands to assess planning capabilities. Furthermore, the bladder and pancreas, which did not appear in the training set, were considered unseen senses for evaluation. The Success Rate (SR) is the fraction of successfully completed instructions. As Table III indicates, USPilot achieved a 99% success rate for single API executions. The fine-tuned adapter can comprehend the instructions, achieving a 95% success rate for direct commands and 78.5% for symptom-based descriptions involving seen body parts. For unseen body parts (bladder and pancreas), USPilot leveraged the LLM’s generalization ability to achieve a 71.4% success rate for direct commands. However, it achieved only a 25.7% success rate for symptom-based descriptions due to the absence of specific knowledge about these body parts in the training data. Instead of effectively generalizing to entirely unseen body parts, it exhibits knowledge transfer between related body parts, such as from the gallbladder to the pancreas or from the kidney to the bladder, which results in lower success rates.
@@
Type Num Accuracy
QA 9167 100%
Robot 5418 97%
@@
IV-C USPilot in the real world ultrasound senario
@@
(b) The graph structure of USPilot with 21 API vertex and 24 edges.
Figure 6: USPilot in a Real World Ultrasound Senario
As shown in Fig. 6(a), the physical setup consists of an ultrasound system (Pioneer H20, Angell, China) equipped with a linear probe and a convex probe mounted on a 6-axis collaborative robotic arm (RM65, Realman, China). A probe fast assemble machine is positioned adjacent to the robot system. A 3D camera (Realsense D455, Intel, US) mounted on the robot platform provides anatomical structure positioning and patient orientation information, serving as part of the system’s environmental feedback. With the patient anatomy structure, the system determines the initial scanning position for different targets. A 6-axis force sensor (M3815D, SRI, China) is integrated into the robot’s end-effector to ensure patient safety and provide orientation feedback for the system. There are 21 APIs and 24 edges defined in the USPilot system, as depicted in the graph structure in Fig 6(b). The system employs PicoVoice as a wake-word listener for user-machine interaction to ensure accessibility and ease of use. The user can activate the system by calling ’Doctor’ or ’Hey, Doctor’. Upon activation, the user’s speech is processed by the VOSK speech recognition toolkit and transcribed into text to USPilot. The system can then convert the generated text or pre-defined API text into audio using EdgeTTS, which is played back to the user.
@@
Carotid vessel 10 I got dizzy recently, the doctor said my carotid vessel may blocked. 10/10
@@
In the real-world scenes, the USPilot was evaluated on two targets with 10 testing rounds per target and 5 additional rounds for single API commands. The robot was required to perform automatic scanning based on varying instructions for each target. We employed the success rate for evaluation, defined as the proportion of testing rounds in which robots successfully completed all assigned tasks. In the scanning task, the robot can successfully locate the commanded target and perform the scan with a 100% success rate. However, for the force adjustment task, USPilot achieved only a 60% success rate, consistent with the results in Table III. This limitation is attributed to the semantic interpretation of the “feeling uncomfortable” instruction, which prompts the system to interrupt the scan rather than solely reducing the contact force.
V Discussion & Future work
The results in Table. I show that the proposed LLMEG method with LLaMA3-8b outperforms GNN4TaskPlan with GPT-4turbo-8*222b. As can be seen from Fig. 5, a reduction in word embedding size significantly impacts the performance of the proposed method, preventing it from achieving the same metrics as the LLaMA3-8b model. Additionally, as the size of the LLMs decreases, the performance of the proposed method declines substantially, indicating that the quality of steps generated by LLMs strongly influences the GNN classification model. Even with a good API selection, the order of the API calls heavily depends on the LLM’s ability to plan subgraphs in the proposed method. Therefore, future research should focus on API selection and optimizing the planning of API call sequences.
As discussed in Section IV-B, the proposed USPilot system provides the potential for an autonomous robotic ultrasound platform. The system effectively distinguishes between user inputs as either questions or robot-executable commands. Nevertheless, the fine-tuned adapter in robotic ultrasound manipulation does provide good performance in executing direct commands and interpreting symptom descriptions. However, the dataset is still not large enough to ensure reliable and stable performance in such a medical scenario. Expanding the dataset with more diverse language instructions would be essential for further fine-tuning.
In section IV-C, we demonstrate the USPilot with the physical setup to perform real ultrasound scans on the human body. The current robotic ultrasound system can achieve unsupervised scans on two simple organs. However, the physical setup still could not provide enough function for medical diagnosis and could not cover as many body parts as the simulated data. Consequently, developing additional applications to address more complex medical scenarios is imperative. Moreover, the current proposed system lacks of the ability of replanning, which should also be discussed in future works.
VI Conclusion

### v2

@@ III-B Comparison experiment for LLMEG
Ours LLaMA3 8b 97.29 82.87 78.64 89.37 75.35 60.8 91.09 74.94 59.6
@@
In the compression experiment, we employed two alternative methods: LLM+LLM: This method involves collaboration between two LLMs. Initially, the text descriptions of all the tools and the user request are provided to the first LLM, which generates the necessary substeps to complete the task. Subsequently, the generated substeps are combined with the tool descriptions to form another prompt for the second LLM, which selects the tools to be utilized. The selected tools are considered as vertexes, while two adjacent vertexes can be regarded as edges. GNN4Task[21]: The GNN4Task generates subtasks using LLMs and employs a small language model (e5-355M) to embed each subtask. A GNN processes each embedded subtask to identify vertexes.
@@
We employed an 8-billion-parameter LLaMA3 model in our proposed method LLMEG and compared it with GNN4Task, which is equipped with one of the powerful LLMs, GPT4-tubo, a mixture of eight 222 billion parameters experts. As shown in Table I , the proposed LLMEG achieves superior performance across all three public datasets. On the Dailylife dataset, both methods perform well due to the inherent ability of LLMs to understand daily service descriptions. On the Multimedia dataset, our method slightly outperforms GNN4Task. On the more professional HuggingFace dataset, it significantly enhances vertex selection and edge prediction. We also compared our method with a variant using DFS instead of the LLM-generated subgraph. Subgraph planning achieves improvements on both edge F1 score and total accuracy, especially on Dailylife dataset. This highlights the importance of semantic understanding in generating correct API sequences, even with accurate vertex selection. The largest gain on Dailylife, which contains 39 edges per vertex, showing that semantic matters more as graph complexity increases.
@@
After employing the LLMEG, we conducted experiments using different LLMs with varying parameter sizes to assess planning capabilities. As shown in Fig. 6(a), our proposed method outperforms the original LLM-based method and GNN4Task in both vertex and edge F1 scores by over 40% across all datasets when using LLaMA3-8b and its embeddings. For other LLMs who cannot generate original sentence embeddings, LLMEG abstracts word embedding with 137M parameters small language model (nomic-embed-textv1.5). Our proposed method achieves better performance across all metrics for a similar size Qwen2.5-7b. However, when integrated with smaller LLMs shown in Fig. 6(c), our LLMEG can conduct a high value only in vertex F1 score but struggles with edge F1 scores and accuracy. This is due to smaller LLMs’ limited semantic understanding capabilities, which are insufficient for effective subgraph replanning.
III-C User intent understanding and ultrasound knowledge embedded
@@
The user intent prediction MLP router was trained with 1.2K data randomly split from the question set with an equal ratio of 1:1:1 for 100 epochs. We used the AdamW optimizer with a learning rate of $61e-$ 6 and a batch size of 1. We used a sigmoid as the activation function to classify the user inputs as QA inquiries or ultrasound robot commands. We also deployed a direct fine-tuning method by mixing the QA and robotic execution data to fine-tune the LLM. The fine-tune is on the same 1.2K train set as the MLP with 10 epochs. Adamw optimizer was used with a learning rate of 1e-5, and a batch size of 1. As shown in Table II, the model was evaluated on the rest of the question set. The results demonstrate that the MLP layer can reach a 95% classification accuracy and the direct fine-tune one can only reach 91% accuracy.
@@
Router Finetune
precision recall precision recall
QA 0.9 0.99 0.97 0.97
Robot 1.0 0.98 0.87 0.84
Accuracy 0.98 0.91
@@
In addition, the ultrasound robotic adapter was fine-tuned on the whole instruction set for 100 epochs. The training set and test set are split with a ratio of 9:1. AdamW optimizer was used with a learning rate of $51e-$ 5, and a batch size of 1. Furthermore, the bladder and pancreas, which did not appear in the training set, were considered as unseen scenes for evaluation to demonstrate the LLM’s ability to transform the knowledge to similar scenarios. The Success Rate (SR) is the fraction of successfully completed instructions. As Table III indicates, USPilot achieved a 99% success rate for single API executions. The fine-tuned adapter can comprehend the instructions, achieving a 95% success rate for direct commands and 78.5% for symptom-based descriptions involving seen body parts. For unseen body parts (bladder and pancreas), USPilot leveraged the LLM’s generalization ability to achieve a 71.4% success rate for direct commands. However, it achieved only a 25.7% success rate for symptom-based descriptions due to the absence of specific knowledge about these body parts in the training data. Instead of effectively generalizing to entirely unseen body parts, it exhibited knowledge transfer between related body parts, such as from the gallbladder to the pancreas or from the kidney to the bladder, which resulted in lower success rates.
III-D USPilot in the real world ultrasound scenario
@@
(b) The graph structure of USPilot.
Figure 7: USPilot in a Real World Ultrasound Scenario
As shown in Fig. 7(a), the physical setup consists of an ultrasound system (Pioneer H20, Angell, China) equipped with a linear probe and a convex probe mounted on a 6-axis collaborative robotic arm (RM65, Realman, China). A probe fast assemble machine is positioned adjacent to the robot system. A 3D camera (Realsense D455, Intel, US) mounted on the robot platform provides anatomical structure positioning and patient orientation information, serving as part of the system’s environmental feedback. With the patient anatomy structure, the system determines the initial scanning position for different targets. A 6-axis force sensor (M3815D, SRI, China) is integrated into the robot’s end-effector to ensure patient safety and provide orientation feedback for the system. There are 21 APIs and 24 edges defined in the USPilot system, as depicted in the graph structure in Fig. 7(b). The system employs a wake-word listener for human-machine interaction to ensure accessibility and ease of use. The user can activate the system by calling “Doctor” or “Hey, Doctor”. Upon activation, the user’s speech is processed and transcribed into text to USPilot. The system can then convert the generated text or pre-defined API text into audio, which is played back to the user.
The computational requirements of the proposed system are influenced by the inference latency, which largely depends on the content and length of the user request. For non-quantized models operating in FP16 precision, LLaMA3-8b requires approximately 20 GB of GPU memory. In our real-world ultrasound scenario experiment, the robot was controlled by a local PC, while the model ran on a local server equipped with a 24 GB Nvidia 4090 GPU. The total latency consists of three main components: 1) The router and request decomposition process requires 0.029 seconds per inference token. On average, this results in a latency of 0.6 seconds per single-step request on our dataset. 2) Using LLaMA3-8b as the backbone, the graph generation process takes an average of 0.13 seconds to produce the final result. 3) Communication between the PC and the local server is handled via Redis, introducing an average latency of 0.47 seconds per interaction.
@@
Carotid vessel 10 I got dizzy recently, the doctor said my carotid vessel may be blocked. 10/10
@@
Figure 8: The USPilot force result with instructions to reduce force
@@
In the real-world scenes, the USPilot was evaluated on two body targets with 10 testing rounds per target and 5 additional rounds for single API commands in Table III. The robot was required to perform automatic scanning based on varying instructions for each target. We used the success rate as the evaluation metric, defined as the proportion of test rounds in which the robots successfully completed all assigned tasks. In the scanning task, the robot can successfully locate the commanded target and perform the scan with a 100% success rate. For the force adjustment task, USPilot achieved a success rate of only 60%, as shown in Fig. 8 , which records the contact force during the experiment. The two failure cases are due to a misunderstanding of the word “painful” of the LLM to interrupt the scan. The time costs were analyzed across three stages: A-B represents the user’s speaking period, B-C corresponds to the speech-to-text transcription time, which ranged from 0.29s to 0.83s, and C-D denotes the LLM and graph processing time, spanning approximately 0.8s to 1.4s. The initial contact force was set to 5 N, with a root mean square error (RMSE) ranging from 0.464 N to 1.102 N. After processing the user’s instructions, the expected force was reduced to 3 N, resulting in an RMSE of 0.446 N to 0.731 N across experiments 1 through 3.
IV Discussion & Future work
The results in Table I demonstrate that the proposed LLMEG method with LLaMA3-8b surpasses GNN4TaskPlan with GPT-4turbo-8*222b. As shown in Fig. 6, reducing word embedding size significantly degrades the performance of the proposed method, preventing it from matching the metrics of the LLaMA3-8b model. Moreover, smaller LLM sizes lead to substantial performance declines, highlighting the strong influence of LLM-generated steps on the GNN classification model. The experiments also reveal that, even with optimal API selection, the order of API calls is heavily dependent on the LLM’s planning capability. As discussed in Section III-C, the proposed USPilot system offers potential for an autonomous robotic ultrasound platform by effectively distinguishing user inputs as either questions or robot-executable commands. While the fine-tuned adapter performs well in executing direct commands and interpreting symptom descriptions, the dataset remains too limited to ensure reliable performance in medical scenarios. Expanding the dataset with more diverse language instructions will be critical for further fine-tuning. In section III-D, we demonstrated the USPilot with the physical setup to perform real ultrasound scans on the human body.
To improve the stability and accuracy, the proposed LLMEG still relies on strong prior knowledge of the robotic system, which limits adaptability. A more generalized approach of using two collaborating LLMs was also evaluated in Section III-B , which requires no extra prior knowledge but performed the worst as shown in Fig. 6 . This sort of trade-off is necessary under our fixed self-built robotic system and protocol-driven nature of ultrasound scanning, where high generalization is typically unnecessary. Learning on priors is reasonable for stability and accuracy, but exploring to reduce training cost remains important.
The proposed method focuses on planning ultrasound robotic procedures, introducing a novel human-robot interaction framework for an unmanned ultrasound system. However, several areas warrant further exploration. Firstly, while the system effectively serves as a planner, arranging robotic tool sequences based on diverse user inputs and acting as a “virtual doctor” to answer questions via a semantic router, it heavily relies on predefined APIs. These APIs dictate the rule-based control logic, limiting the system’s adaptability and scalability. A key challenge during development was designing an API structure that balanced diverse user instructions with compatibility to rule-based mechanisms. With advancements in multimodal vision-language-action models in collaborative robotics, future work should integrate real-time probe control into the framework. Such models could process not only language inputs but also image and force data, enabling adaptive, context-aware manipulation during ultrasound procedures. Secondly, the graph-based structure in the proposed method uses undirected graph features due to the nature of the GNN. This limits its ability to handle scenarios requiring cyclic graph structures, which are important for dynamic replanning and learning from errors. Incorporating directed graph features or alternative learning mechanisms will be essential to enhance the system’s robustness in managing unexpected scenarios or user modifications. While the current method provides a front-end for an automated sonographer, developing a more flexible and intelligent back-end remains a critical challenge for future research.
V Conclusion